# Stepwise Regression and Pre-trained Edge for Robust Stereo Matching

### v3

@@ Abstract
Due to the difficulty in obtaining real samples and ground truth, the generalization performance and the fine-tuned performance are critical for the feasibility of stereo matching methods in real-world applications. However, the diverse datasets exhibit substantial discrepancies in disparity distribution and density, thus presenting a formidable challenge to the generalization and fine-tuning of the model. In this paper, we propose a novel stereo matching method, called SR-Stereo, which mitigates the distributional discrepancies across different datasets by predicting the disparity clips and uses a loss weight related to the regression objective scale to improve the accuracy of the disparity clips. Moreover, this stepwise regression architecture can be easily extended to existing iteration-based methods to improve the performance without changing the structure. In addition, to mitigate the edge blurring of the fine-tuned model on sparse ground truth, we propose Domain Adaptation Based on Pre-trained Edge (DAPE). Specifically, we use the predicted disparity and RGB image to estimate the edge map of the target domain image. The edge map is filtered to generate edge map background pseudo-labels, which together with the sparse ground truth disparity on the target domain are used as a supervision to jointly fine-tune the pre-trained stereo matching model. These proposed methods are extensively evaluated on SceneFlow, KITTI, Middbury 2014 and ETH3D. The SR-Stereo achieves competitive disparity estimation performance and state-of-the-art cross-domain generalisation performance. Meanwhile, the proposed DAPE significantly improves the disparity estimation performance of fine-tuned models, especially in the textureless and detail regions.
@@
Stereo, robust stereo matching, stepwise regression, domain adaptation
@@ I Introduction
Depth information is key to computer vision and graphics research in the real world. Accurate depth estimation is vital for fields such as autonomous driving, robot navigation and action recognition. Common depth estimation methods include structured light [1], stereo matching [2, 3, 4], radar [5], etc. The structured light method projects a coded pattern onto the object surface and estimates the depth by observing the distortion of the pattern imaged on the object surface. Due to the need to project multiple patterns consecutively, structured light methods are generally only suitable for static indoor scenes. The radar obtains a sparse depth map by transmitting pulses and receiving echoes. Millimeter-wave radar has a long detection range but fuzzy spatial resolution, while laser radar has a high spatial resolution but is expensive. In contrast, stereo matching calculates the depth map by estimating the horizontal displacement map (i.e., the disparity) of the pixels between the corrected left and right image pairs. By utilizing only two cameras, stereo matching is able to provide dense depth maps in different scenes, and is therefore considered a cost-effective and widely applicable method for depth estimation.
@@
Figure 1: Domain-adaptive visualization on KITTI. All methods are trained on SceneFlow and fine-tuned on KITTI. During inference, all methods run 15 disparity updates. Our proposed SR-Stereo achieves more accurate performance in the edge region. Meanwhile, as a generalized fine-tuning framework, our proposed DAPE effectively improves the disparity estimation performance of existing methods fine-tuned with sparse ground truth.
@@
Many learning-based stereo methods [12, 6, 7, 8, 9, 11, 10] have achieved encouraging success. Some works [12, 13, 14] focus on designing the 4D cost volume, which characterizes the cost of pixel matching between left and right image pairs. They use 3D convolutional networks to aggregate and regularize the entire cost volume, and regress it to obtain the disparity. However, the large number of 3D convolutions leads to large computational and memory cost, which limits the applicability of such methods. Recently, iteration-based methods [16, 15, 17, 22, 20, 18, 21, 19] have shown great potential in terms of both accuracy and real-time performance, and thus have become the mainstay of current research. Iterative-based methods index the similarity features within current disparity neighborhood from the cost volume and then use it to update the disparity. This approach avoids the expensive cost of aggregating and regularizing cost volume and achieves a balance between performance and efficiency by controlling the number of iterations. In addition, some state-of-the-art iteration-based methods [17, 18] utilize a lightweight 3D convolutional network to regress the coarse initial disparity, thus further accelerating the iteration efficiency.
@@
However, stereo methods usually suffer from the problems of few real samples and difficult access to ground truth in practical applications. Existing methods [26, 25, 27, 24, 23] are generally pre-trained on a virtual dataset (e.g., SceneFlow [29]) and fine-tuned or directly inferred on real datasets, which imposes higher requirement on the generalization performance. In addition, there are significant discrepancies in disparity distribution and density among different datasets due to variations in acquisition equipment and environment. The Middlebury 2014 [30], which uses hand labeling to obtain dense ground truth, has a maximum disparity value of more than 250 pixels. The KITTI dataset  [32, 31], which uses radar and careful calibration to obtain sparse ground truth, has all of its disparities within 100 pixels. These discrepancies make it difficult for a method that performs well on one dataset to directly achieve the same performance on other datasets. Therefore, we need to improve the method so that it has good performance in estimating disparity across domains.
In this paper, we claim that by designing a regression objective with a similar distribution across different datasets, excellent performance can be consistently achieved using the same method across various datasets. Therefore, a disparity stepwise regression (DSR) is proposed to study the disparity estimation in robust stereo matching. Specifically, the regression objective in the DSR is multiple disparity clips, which have the same range across different datasets. By summing the multiple disparity clips, the DSR enables accurate and robust disparity estimation. Meanwhile, a loss weight related to the regression objective scale is proposed to improve the accuracy of disparity clips because it effectively mitigates the long-tailed distribution problem [33, 34] of the disparity clips. Based on these two simple but powerful methods, we design a novel architecture, Stepwise Regression Stereo (SR-Stereo), for robust disparity estimation.
@@
Furthermore, we observe that models fine-tuned with sparse ground truth suffer from severe edge blurring, i.e., edge disparity errors. As shown in Figure 1, even the state-of-the-art models [16, 17] in the KITTI online leaderboard fail to accurately estimate the edge disparity. This further limits the effectiveness of stereo methods in real-world applications, as the vast majority of efficient labeling methods (e.g., radar) can only obtain sparse ground truth. The most direct way to mitigate this problem is to introduce edge-dependent supervision, but current edge feature extraction methods [35, 36, 37, 38, 39, 40] have difficulty in obtaining accurate and concise edges as pseudo-label due to the overly sparse disparity ground truth and the interference of object textures in RGB images.
@@
In this paper, a Domain Adaptation Based on Pre-trained Edge (DAPE) is proposed to address the issue of blurring in edge details. The two key observations behind this approach are: 1) disparity has more concise and reliable information about object edges compared to RGB maps. 2) models pre-trained with dense ground truth tend to have accurate edge in predicted disparity on new domains. Therefore, we first propose a lightweight edge estimator that takes both disparity and RGB image as inputs to generate a dense edge map. This estimator, along with SR-Stereo, is pre-trained on a virtual dataset. We utilize the pre-trained SR-Stereo to generate the disparity map for the target domain, which is further fed into the edge estimator to predict the edge map. Then, the background pixels in the edge map are retained and utilized as edge pseudo-label, while the foreground pixels are filtered out. Finally, we use the edge pseudo-label as additional supervision during the model fine-tuning process, which effectively improves the model’s disparity estimation performance on edge details.
@@
  • We propose a novel architecture, called SR-Stereo, which predicts disparity clips to alleviate distribution discrepancies across different datasets and uses a loss weight related to the regression objective scale to improve the accuracy of the disparity clips. Experiments on RAFT-Stereo and IGEV-Stereo also demonstrate that the proposed stepwise regression architecture has better generalization performance than iteration-based methods.
  • We propose a novel fine-tuning framework for real datasets, named DAPE. By employing generated edge pseudo-label to supervise the model fine-tuning process, our proposed method effectively improves the disparity estimation performance of models fine-tuned with sparse ground truth.
  • We perform experiments on SceneFlow, KITTI, Middlebury 2014 and ETH3D. Compared to existing methods, the proposed SR-Stereo achieves the state-of-the-art cross-domain generalization performance across multiple datasets and obtains competitive results on the SceneFlow and KITTI benchmark. Meanwhile, experimental results on multiple realistic datasets show that the proposed DAPE significantly improves the disparity estimation performance of the fine-tuned model, especially in untextured and detailed regions.
@@ II-A Iterative-based Stereo Matching
Compared to cost aggregation-based methods [41, 42, 43], iteration-based methods achieve significant improvement in both efficiency and accuracy. RAFT-Stereo [44, 16] is the first iteration-based method that constructs cost volume by calculating the dot product of features with same height. At the same time, it introduces multi-level GRU (Gated Recurrent Unit) unit [45] which updates the disparity by indexing the features of cost volume. Based on RAFT-Stereo, DLNR [22] uses decoupled LSTM [46] (Long Short-Term Memory) instead of GRU to retain more high-frequency information during the iteration process, and designs a normalized refinement module to capture more detailed information at full resolution. The refinement of high-frequency information effectively improves the performance of disparity estimation. Further, CREStereo [15] applies the iterative process to different resolutions of disparity. It uses a deformable search window to more accurately index the features of cost volume, and proposes a hierarchical refinement network to update the disparity at different resolutions in a coarse-to-fine order. In recent research, cost aggregation based on 3D convolutional networks is revisited. For instance, IGEV-Stereo [17] constructs a geometric encoding volume by combining Group-wise correlation and All-pairs Correlation, and uses a lightweight 3D regularization network to regress the initial disparity, which effectively improves the iterative efficiency and accuracy.
Our proposed SR-Stereo is similar in process to the iteration-based methods, but has a different objective. The iteration-based methods focus on how to efficiently converge the disparity, and thus all of its update units target the ground truth disparity. Instead, the proposed SR-Stereo focuses on improving the generalization performance by mitigating the discrepancies in the disparity distribution among different domains. To this end, the update units of SR-Stereo are all individually assigned different objectives, i.e., range-constrained disparity clips.
II-B Robust Stereo Matching
Since real samples and ground truth disparity are difficult to obtain for practical applications, some works [26, 27, 28] focus on cross-domain generalization, which aims to ensure good performance of the model on unseen datasets. CFNet [26] and UCFNet [27] narrow the domain discrepancy by cascaded cost volume and multi-stage refinement of disparity. They also adaptively adjust the search space of disparity at each stage by uncertainty estimation. DSMNet [28] employs domain normalization and a trainable structure-preserving graph-based filter to extract robust geometric feature to improve the cross-domain generalization performance. These methods enhance the generalization performance by improving the network structure. In contrast, our proposed DSR designs a new regression objective with similar distribution across different datasets to reduce the generalization difficulty.
Figure 2: The overall framework of the proposed DAPE. First, a robust stereo model SR-Stereo and a lightweight edge estimator are pre-trained on a large synthetic dataset with dense ground truth. Then, we use the pre-trained SR-Stereo and edge estimator to generate the edge map of target domain, where the background pixels (i.e., non-edge region pixels) are used as edge pseudo-label. Finally, we jointly fine-tune the pre-trained SR-Stereo using the edge pseudo-label and sparse ground truth disparity.
@@ II-B Robust Stereo Matching
In addition, making good use of the few real samples is important for domain adaptation of the model. Some works [17, 16, 18] have demonstrated that even using only a few samples, the performance of pre-trained models in a new domain can be significantly improved. However, most existing methods simply employ direct fine-tuning of the model using the ground truth disparity of the new domain, without considering the impact of ground truth density on domain adaptation. For instance, models fine-tuned with sparse ground truth often suffer from severe edge blurring. To address this problem, we propose a domain adaptation based on pre-trained edge which utilizes edge pseudo-label from the target domain to preserve the model’s disparity estimation capability for edge details.
@@ III METHODOLOGY
III-A Framework Overview
In this paper, we propose a novel stereo matching architecture, SR-Stereo, which achieves robust disparity estimation by stepwise regression of the disparity. In addition, we provide a general and reliable framework, DAPE, for model fine-tuning on sparse ground truth by digging into the disparity-based edge estimation. As shown in Figure 2, the proposed DAPE consists of three steps:
  1. 1.
    Firstly, a robust stereo model and a lightweight edge estimator are pre-trained on a large synthetic dataset with dense ground truth. Specifically, a stepwise regression stereo (SR-Stereo) is proposed to mitigate the disparity distribution discrepancies between different datasets. Meanwhile, the disparities predicted by SR-Stereo are fed jointly with the RGB images into a lightweight network to learn the edge information.
  2. 2.
    Then, the pre-trained stereo matching model and the edge estimator are used to directly infer on the target domain and generate corresponding edge maps. To mitigate the negative impact of disparity errors in reflective regions, we propose to use only the relatively dense background pixels (i.e., non-edge region pixels) from the edge maps as pseudo-label.
  3. 3.
    After generating the edge pseudo-label of the target domain, we use it and the sparse ground truth disparity to jointly fine-tune the pre-trained stereo matching model. Experimental results show that the incorporation of edge pseudo-label supervision significantly improves the disparity estimation performance of the fine-tuned model.
The following subsections describe the proposed SR-Stereo and DAPE in more detail.
III-B Stepwise Regression Stereo

### v4

# SR-Stereo & DAPE: Stepwise Regression and Pre-trained Edges for Practical Stereo Matching

@@ Abstract
Due to the difficulty in obtaining real samples and ground truth, the generalization performance and domain adaptation performance are critical for the feasibility of stereo matching methods in practical applications. However, there are significant distributional discrepancies among different domains, which pose challenges for generalization and domain adaptation of the model. Inspired by the iteration-based methods, we propose a novel stepwise regression architecture. This architecture regresses the disparity error through multiple range-controlled clips, which effectively overcomes domain discrepancies. We implement this architecture based on the iterative-based methods , and refer to this new stereo method as SR-Stereo. Specifically, a new stepwise regression unit is proposed to replace the original update unit in order to control the range of output. Meanwhile, a regression objective segment is proposed to set the supervision individually for each stepwise regression unit. In addition, to enhance the edge awareness of models adapting new domains with sparse ground truth, we propose Domain Adaptation based on Pre-trained Edges (DAPE). In DAPE, a pre-trained stereo model and an edge estimator are used to estimate the edge maps of the target domain images, which along with the sparse ground truth disparity are used to fine-tune the stereo model. The proposed SR-Stereo and DAPE are extensively evaluated on SceneFlow, KITTI, Middbury 2014 and ETH3D. Compared with the SOTA methods and generalized methods, the proposed SR-Stereo achieves competitive in-domain and cross-domain performances. Meanwhile, the proposed DAPE significantly improves the performance of the fine-tuned model, especially in the texture-less and detailed regions. The code is available at https://github.com/zhuxing0/SR-Stereov1-DAPE.
@@
Stereo matching, generalization performance, stepwise regression, domain adaptation.
@@ I Introduction
Depth information is key to computer vision and graphics research in the real world. Accurate depth estimation is vital for fields such as autonomous driving, robot navigation and action recognition. Common depth estimation methods include structured light [1], stereo matching [2, 4, 5, 6, 7], and radar [3], etc. The structured light method projects a coded pattern onto the object surface and estimates the depth by observing the distortion of the pattern imaged on the object surface. Due to the need to project multiple patterns consecutively, structured light methods are generally only suitable for static indoor scenes. The radar obtains a sparse depth map by transmitting pulses and receiving echoes. Millimeter-wave radar has a long detection range but fuzzy spatial resolution, while laser radar has a high spatial resolution but is expensive. In contrast, stereo matching calculates the depth map by estimating the horizontal displacement map (i.e., the disparity) of the pixels between the corrected left and right image pairs. By utilizing only two cameras, stereo matching is able to provide dense depth maps in different scenes, and is therefore considered a cost-effective and widely applicable method for depth estimation.
@@
Figure 1: Comparison of reconstructed point clouds on KITTI. All methods are trained on SceneFlow and fine-tuned on KITTI. During inference, all methods run 15 disparity updates. Our SR-Stereo performs better in the detailed regions. In addition, the proposed fine-tuning framework DAPE effectively improves the performance of existing methods fine-tuned on sparse ground truth.
@@
Many learning-based stereo methods [5, 4, 6, 7, 9, 8, 10, 13, 11, 12] have achieved encouraging success. Some works [5, 6, 7] focus on designing the 4D cost volume, which characterizes the cost of pixel matching between left and right image pairs. They use 3D convolutional networks to aggregate and regularize the entire cost volume, and regress it to obtain the disparity. Recently, iteration-based methods [9, 8, 10, 13, 11, 12] have shown great potential in terms of both accuracy and real-time performance, and thus have become the mainstay of current research. The iteration-based methods use similarity features from the cost volume to predict the disparity error for optimization of the disparity. This approach avoids the expensive cost of aggregating and regularizing cost volume and achieves a balance between performance and efficiency by controlling the number of iterations.
@@
However, real training samples and ground truth is usually difficult to obtain for stereo methods in practical applications. Existing methods [14, 23, 24, 25] are generally pre-trained on a large synthetic dataset (e.g., SceneFlow [16]) and then fine-tuned or directly inferred on real datasets, which makes higher demands on the generalization performance and domain adaptation performance. In addition, there are significant distributional discrepancies among different datasets due to differences in scene and resolution. For example, the maximum disparity in the Middlebury [17] exceeds 250 pixels, while all the disparities in the KITTI [19, 18] are within 100 pixels. These discrepancies make it difficult for a method that performs well on one dataset to directly achieve the same performance on other datasets. Therefore, a new architecture needs to be designed to overcome the domain discrepancies to achieve better robustness.
Our architecture is inspired by iteration-based methods. Iteration-based methods aim to correct disparity by directly predicting the current disparity error using update units. However, influenced by the disparity distribution, the disparity errors tend to vary across domains, which limits the generalization performance of iteration-based methods. Additionally, tens of updates are usually required to obtain accurate results. Based on these observations, we propose a novel stepwise regression architecture. In this architecture, we split the disparity error into multiple segments with fixed ranges and use multiple update units to predict them separately. This architecture has the following advantages: 1) the segments are range-controlled and domain-independent, and thus easy to generalize. 2) it still requires only tens of updates, and does not introduce additional computational and time costs.
We implement this architecture on the basis of iteration-based methods, and refer to this new stereo method as SR-Stereo. Specifically, we propose a stepwise regression unit to replace the original update unit. The output of the stepwise regression unit is range-controlled rather than unconstrained as in the original update unit. Then, we propose a regression objective segment, which sets a separate regression objective for each stepwise regression unit based on the splitting result of the disparity error. Further, we introduce a Disparity Clip-Balanced Weight to improve the accuracy of the predicted segments in the stepwise regression units.
@@
Furthermore, we observe that models fine-tuned with sparse ground truth suffer from severe edge blurring, i.e., edge disparity errors. As shown in Figure 1, even the state-of-the-art models [10, 11] in the KITTI online leaderboard fail to accurately estimate the edge disparity. This further limits the effectiveness of stereo methods in real-world applications, as the vast majority of efficient labeling methods (e.g., radar) can only obtain sparse ground truth. The most direct way to mitigate this problem is to introduce edge-dependent supervision, but current edge feature extraction methods [20, 21, 22] have difficulty in obtaining accurate and concise edge information due to the overly sparse disparity ground truth and the interference of object textures in RGB images.
@@
In this paper, we propose a novel model fine-tuning framework to address the problem of blurring in edge details, which we name Domain Adaptation based on Pre-trained Edge (DAPE). The two key observations behind this framework are: 1) in contrast to RGB images, in disparity, the texture of the object is weakened while concise edge information is highlighted. 2) for the model pre-trained with dense ground truth, its predicted disparity on new domains tends to have accurate edge contours. Therefore, we first propose a lightweight edge estimator that takes both disparity and RGB image as inputs to generate a dense edge map. This estimator, along with SR-Stereo, is pre-trained on a virtual dataset. We utilize the pre-trained SR-Stereo to generate the disparity map for the target domain, which is further fed into the edge estimator to predict the edge map. Then, the background pixels in the edge map are retained, while the foreground pixels are filtered out. Following UCFNet [14], these pixels are defined as pseudo-labels. Finally, we use the edge pseudo-labels as additional supervision during the model fine-tuning process, which effectively improves the model’s disparity estimation performance on edge details.
@@
  • We propose a novel stereo method called SR-Stereo. It splits the disparity error into multiple segments for regression, which effectively overcomes the domain discrepancies (i.e., differences in disparity distribution). More importantly, we propose a stepwise regression architecture, which provides a new implementation idea for generalized stereo matching.
  • We propose a novel fine-tuning framework for real datasets, named DAPE. By employing generated edge pseudo-label to supervise the model fine-tuning process, the proposed DAPE effectively improves the disparity estimation performance of models fine-tuned with sparse ground truth.
  • We perform extensive experiments on SceneFlow, KITTI, Middlebury and ETH3D. Compared to the state-of-the-art methods, our SR-Stereo achieves competitive performance on several benchmarks. Meanwhile, compared to the majority of generalized stereo methods, our SR-Stereo achieves the best performance on Middlebury and ETH3D. The generalization performance of our method on KITTI is also very competitive among the methods of the last two years. In addition, experimental results on multiple realistic datasets show that DAPE significantly improves the performance of the fine-tuned model, especially in texture-less and detailed regions.
@@ II-A Iterative-based Stereo Matching
Compared to cost aggregation-based methods, iteration-based methods [23, 24] achieve significant improvement in both efficiency and accuracy. RAFT-Stereo [9] is the first iteration-based method that innovatively introduces multi-level GRU [27] (Gated Recurrent Unit) to update disparities in stereo matching. Based on RAFT-Stereo, DLNR [13] replaces GRU with decoupled LSTM [28] (Long Short-Term Memory) to retain more high-frequency information during the iterative process, and designs a normalized refinement module to capture more detailed information at full resolution. Furthermore, CREStereo [8] applies the iterative process to different resolutions of disparity and proposes a hierarchical refinement network to update disparities in a coarse-to-fine manner. IGEV-Stereo [10] constructs a combined geometric encoding volume by combining Geometric Encoding Volume and All-Pairs Correlation, and uses a lightweight 3D regularization network to regress a rough initial disparity, which effectively improves the iterative efficiency and accuracy. Selective-Stereo [11] further enhances performance by extracting disparity information of different frequencies using GRUs with varying kernel sizes. Overall, these methods aim to directly regress the entire disparity error to achieve fast disparity optimization. In contrast, our SR-Stereo does not insist on directly regressing the disparity error, but instead splits it into multiple range-controlled and domain-independent segments to overcome the distributional discrepancies among different domains.
II-B Generalized Stereo Matching and Domain Adaptation
Several works have focused on cross-domain generalization and domain adaptation. UCFNet [14] narrows domain discrepancies by refining disparities in stages and adaptively adjusts the disparity search space at each stage through uncertainty estimation. Additionally, it proposes a fine-tuning framework based on pseudo-labels of target domain disparities to enhance the model’s domain adaptation performance. HVT [39] enriches the distribution of training data by transforming a synthetic dataset at three levels (global, local, and pixel), thereby improving the model’s generalization ability. GraftNet [37] incorporates broad-spectrum features from a large-scale dataset into stereo matching to enhance the model’s robustness to image styles. Rao et al. [40] enhance the learning of structural information by integrating stereo matching and image reconstruction tasks, thereby improving the model’s generalization performance. Zhang et al. [38] explicitly constrain the feature consistency of matching pixel pairs by introducing a stereo contrastive feature loss function. DKT [36] introduces a teacher-student architecture, which analyzes the differences between ground truth labels and pseudo-labels to enhance the model’s robustness. In contrast, our SR-Stereo focuses on constructing domain-independent regression objectives, specifically by predicting range-controlled disparity error segments to achieve more robust performance.
@@ II-B Generalized Stereo Matching and Domain Adaptation
In addition, making good use of the few real samples is important for domain adaptation of the model. Some works [10, 9, 11] have demonstrated that even using only a few samples, the performance of pre-trained models in a new domain can be significantly improved. However, these methods simply fine-tune the model using the ground truth disparity of the new domain, without considering the impact of ground truth density on domain adaptation. For instance, models fine-tuned with sparse ground truth often suffer from severe edge blurring. EdgeStereo [41] mitigates this issue by integrating stereo matching and edge detection tasks. Our DAPE provides a simpler solution by using a pre-trained stereo model and edge estimator to offer additional edge supervision for model fine-tuning. In contrast to EdgeStereo, the stereo model and edge estimator in DAPE are completely decoupled, thus avoiding unnecessary overhead during the inference.
@@ III METHODOLOGY
III-A Overview
In this chapter, we provide a novel and complete solution for cross-domain generalization and domain adaptation for stereo methods. Firstly, we propose a novel stereo method, SR-Stereo, which overcomes the domain discrepancies by regressing multiple range-controlled disparities. Then, we propose a general and reliable framework, DAPE, for domain adaptation on sparse ground truth.
Objectively, SR-Stereo and DAPE are independent, i.e., they can be used independently to solve specific problems. In Section III-B, we describe the principle of existing iteration-based methods and highlight the essential improvements introduced by the proposed SR-Stereo. In Section III-C and Section III-D, we describe the proposed SR-Stereo and DAPE in detail.
Figure 2: The disparity update process of iteration-based methods. In this process, the update unit outputs the residual disparity to update the current disparity. The ///// / means stop gradient.
III-B Preliminaries
Distinct from the cost aggregation-based methods, the iteration-based methods employ update units to predict and optimize the disparity. Ignoring steps such as upsampling, the disparity update can be roughly summarized by the process in Figure 2. Obviously, the whole process can be formulated as:
$d_{k}=d_{k-1}^{*}+\Delta d_{k}$ (1)
where $d_{k}$ denotes the predicted disparity after $k$ iterations, $\Delta d_{k}$ denotes the residual disparity in the $k$-th iteration, and * denotes no gradient. Then, during training, all the predicted disparities are supervised:
$Loss=\sum_{k=1}^{N}\gamma_{k}||d_{gt}-d_{k}||$ (2)
where $Loss$ denotes the training loss, $d_{gt}$ denotes the ground truth disparity, $\gamma_{k}$ is the loss weight (hyperparameter) and ||.||||.||| | . | | denotes common loss functions (e.g., L1, SmoothL1, etc.). Considering Eq 1, the supervision in Eq 2 can be equated:
$Loss=\sum_{k=1}^{N}\gamma_{k}||d_{k-1}^{error}-\Delta d_{k}||$ (3)
in which
$d_{k-1}^{error}=d_{gt}-d_{k-1}^{*}$ (4)